OpenStudy (anonymous):

How many kg of coal would be required to produce 1,000 BTU?

thomaster (thomaster):

1 kg coal = 15,000-27,000 kilojoule = 15-27 million joule 1 BTU = 1054–1060 joule 1000 BTU = 1 054 000 - 1 060 000 joule now divide 1054000 by 15000000 = 0.07 kg = 70 g then divide 1060000 by 27000000 = 0.04 kg = 40 g So 40-70 grams of coal would be required to produce 1000 BTU